OpenStudy (anonymous):

find dy/dx for y= (x^x.(cosx sinx)^1/2)/(1+ln(2x))

OpenStudy (rogue):

Is this what you have to differentiate?\[y = \frac { x^x \sqrt {\cos x \sin x} }{1 + \ln (2x)}\]

OpenStudy (anonymous):

yes. haha i hardly can write the equation signs :)